OpenStudy (anonymous):

What is the solution to the equation x – 9 = –4? i think its D A.5 B.–13 C.–36 D.–5

OpenStudy (anonymous):

you add 9 to both sides

OpenStudy (anonymous):

so it is d

OpenStudy (anonymous):

It's not D. Add 9 to both sides. What do you think the answer is now?

OpenStudy (anonymous):

c..?

OpenStudy (anonymous):

(-5) - 9 = (-5) + (-9) = (-14). So it would not be D

OpenStudy (anonymous):

(-36) - 9 = (-36) + (-9) = (-45) So it would not be C

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Think about this. A = 5. Solve for x. 5 - 9 = ?

OpenStudy (anonymous):

im so confused!!

OpenStudy (anonymous):

so it has to be a

OpenStudy (anonymous):

Solve for x. Try all of the answers in place for x and see which one comes out to the answer (-4). ex. a.5 - 9 =? b.(-13) - 9 = (-13) + (-9) =? c.(-36) - 9 = (-36) + (-9) = d.(-5) - 9 = (-5) + (-9) = ^ That is all the problems you have to work out. i have labeled which answer is in place of x. Solve those problems, and see which one leads to (-4)